OpenStudy (anonymous):

I am confused with proving the vector space axiom of c(u+v) = cu + cv for the set V of positive real numbers with addition and scalar multiplication defined as follows: x+y = xy and cx=x^c The proof for that axiom is: c*(x + y) = c*(xy) = (xy)^c = (x^c)(y^c) = (c*x)(c*y) = c*x + c*y I got lost after the (c*x)(c*y) point... How does that translate into the final step? Shouldn't it look like: c^2(xy)?

OpenStudy (anonymous):

Comment: a friend pointed out that I may be confused with scalar/vector... But since c is actually a exponent, you can't turn the expression of (c*x)(c*y) into c*x + c*y using addition function, can you?

OpenStudy (amistre64):

are these 2 separate problems? or is that the conditions which must be met fro x and y

OpenStudy (amistre64):

x+y = xy and cx=x^c x+y-xy = 0 x(1-y)+y = 0 x(1-y)=-y x = y/(y-1) ; y not= 1 x^c = cy/(y-1) hmm

OpenStudy (anonymous):

c is considered a scalar while x and y are vectors

OpenStudy (amistre64):

oh, i wasnt playing with them as vectors was i :)

OpenStudy (amistre64):

then xy means dot product right?

OpenStudy (amistre64):

x.y = x1y1 + x2y2 + ... + xnyn but that wouldnt make sense either since dot product produces a scalar while addition produces another vector

OpenStudy (amistre64):

oh well, i cant make sense out of it, sorry :)

OpenStudy (anonymous):

maybe i am making it too hard heres the website. look at example 5 http://tutorial.math.lamar.edu/Classes/LinAlg/VectorSpaces.aspx

OpenStudy (anonymous):

part g, that is... i am confused with the definition of addition step.... how can we go from (x^c)(y^c) to (x^c) + (y^c)? doesnt that violate the standard addition rule?

OpenStudy (amistre64):

ok, so its a self defined addition and scalar multiplication for the vector space.

OpenStudy (anonymous):

but i am confused about that because it is defined as x+y=xy, not x^c + y^c = (xy)^c

OpenStudy (amistre64):

and your concern is in with g specificaly? right

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

g is the distributive property c(x+y) = cx + cy in general

OpenStudy (anonymous):

that i agree with. but i dont see how its applicable when you have (x^c)(y^c)...... we have exponents there

OpenStudy (amistre64):

is cx defined as x^c?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

then by the definition that they provide; its consistent

OpenStudy (anonymous):

when x^c stands alone, i can see that. but when it is with y^c, i dont see how we can turn it into an expression that looks like addition as shown in part g

OpenStudy (amistre64):

c(x+y) = c(xy); say xy = n cn = n^c (xy)^c x^c y^c x^c is a; y^c = b ab = a+b =x^c + y^c

OpenStudy (amistre64):

these relate to some real positive number; in the scheme provided at the start of example 5

OpenStudy (amistre64):

they are NOT operations that have anything to do with what math is normally defined as

OpenStudy (amistre64):

but we could argue that they are rules of logs

OpenStudy (amistre64):

log(a) + log(b) = log(ab) c log(a) = log a^c

OpenStudy (amistre64):

work it out like that and see if it makes any sense to you :)

OpenStudy (amistre64):

and also, logs only are defined for postive real numbers greater than 0 as well :)

OpenStudy (amistre64):

c ( log(x)+log(y)) = c log(xy) c log(xy) = log(xy)^c = log(x^c y^c) = log(x^c) + log(y^c) log(x^c) + log(y^c) = c log(x) + c log(y)

OpenStudy (amistre64):

remove logs and youve got example 5

OpenStudy (anonymous):

I follow your steps but when you turned x^c to a and y^c to b to perform the addition, I don't buy that. I mean it runs against logic. At least in the way I see it

OpenStudy (anonymous):

Any better way to explain it? Thanks for your patience!

OpenStudy (anonymous):

You also said that they are not standard math operations. How can I truly understand that?

OpenStudy (amistre64):

they mimic the math of logs .....

OpenStudy (anonymous):

:how so?

OpenStudy (anonymous):

c*(x + y) = c*(xy) = (xy)^c = (x^c)(y^c) = (c*x)(c*y) = c*x + c*y The last equal sign is by virtue of your definition of addition